OpenStudy (anonymous):

What is the y-intercept of the graph: 3x – 9y = 18. A. (0, 2) B. (0, -2) C. (-2, 0) D. (0, 6)

OpenStudy (anonymous):

y=(3/9)x-(18/9) equating with y =mx+c where c= intercept of y ans is B